Without graphing, determine whether the system has one solution, no solution, or an infinite number of solutions.
6x - y = 14
x + 3y = 15
thank you

Answer by Nate(3500)

Isolate 'y':
6x - y = 14
-y = -6x + 14
y = 6x - 14
and
x + 3y = 15
3y = -x + 15
y = (-1/3)x + 5
One Solution